In 1854, Mexico sold the United States roughly 30,000 square miles of land known as the Gadsden Purchase for $0.003 per square mile. In 2019, the average value of an acre at this location is $5000. Assuming annually compounded interest, what has been the average annual percentage increase in the value of an acre of this land? (Hints: assume that N, the number of years between 1854 and 2019, is equal to 165; you do not need to consider what day of the year the purchase was sold. Also, be sure to convert between acres and square miles appropriately.) Express your answer in terms of a percentage value, to the nearest one-hundredth percentage point (e.g., if you believe the answer is 1.23%, enter 1.23, not 0.0123).
Answer #1

Total price paid by US in 1854 = 30000 * 0.003 = 90

1 Sq mile = 640 acres

30000 sq miles = 640 * 30000 = 19200000 acres

Value in 2019 = 19200000 * 5000 = 96000000000

F = P *(1+i)^t

96000000000 = 90 * (1+i)^165

(1+i)^165 = 96000000000 / 90 = 1066666667

1+i = 1066666667 ^ (1/165) = 1.134267

i = 0.134267 ~ 13.43%